# Localized Diffusion Models for High Dimensional Distributions Generation

### v1

@@ 3.3.2 A Complete Error Analysis
  1. 1.
    (Boundedness) $p_{0}$ is supported on $[-M,M]^{d}$, and its density is upper and lower bounded by some constants $C_{p},C_{p}^{-1}$ respectively.
  2. 2.
@@
[32] only considers the standard domain $[-1,1]^{d}$. It can be simply extended to $[-M,M]^{d}$ by scaling argument. Denote $p^{M}:=M^{d}p_{0}(M\cdot)$, then $p^{M}$ is supported on $[-1,1]^{d}$ and satisfies the same regularity conditions. Note the scaling only affects the radius $R$ of the Besov space, and does not change the scaling of the sample complexity.
@@
$\mathsf{L}^{j}=\mathcal{O}(\log^{4}n_{j}),\quad\left\|\mathsf{W}^{j}\right\|_{\infty}=\mathcal{O}(n_{j}\log^{6}n_{j}),\quad\mathsf{S}^{j}=\mathcal{O}(n_{j}\log^{8}n_{j}),\quad\mathsf{B}^{j}=n_{j}^{\mathcal{O}(\log\log n_{j})}.$
@@
$\mathrm{e}^{-cr^{*}}+N^{-\frac{2\gamma}{d_{\rm eff}^{*}+2\gamma}}\ll N^{-\frac{2\gamma}{d+2\gamma}}.$
@@
$\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{W}_{2}^{2}(p,\widehat{q}_{T-{\underline{t}}})]\lesssim\mathsf{W}_{2}^{2}(p,p_{\underline{t}})+\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{W}_{2}^{2}(p_{\underline{t}},\widehat{q}_{T-{\underline{t}}})]\lesssim dN^{-k}+\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{KL}(p_{\underline{t}}\|\widehat{q}_{T-{\underline{t}}})].$
@@ 4.1.1 Approximate locality
$P_{t}:=-\nabla^{2}\log p_{t}=(\alpha_{t}^{2}C_{0}+\sigma_{t}^{2}I)^{-1}.$
@@
Figure 1: Top-left: The precision matrix $P_{t}$ at $t=0.1039$, plotted in $\log|P_{t}|$ scale. We can see precise exponential decay of $P_{t}(i,j)$ in $|i-j|$. The rest plots are the localization radius $r_{\rm loc}(t)$ (4.1) under different problem dimension $d$, precision matrix bandwidth $r_{0}$ and condition number $\kappa$. Top-right: $r_{\rm loc}(t)$ with different dimensions. Here $r_{0}=10$ and the condition numbers are similar ($\kappa\approx 193,191,197$). Bottom-left: $r_{\rm loc}(t)$ with different bandwidths. Here $d=1,000$ and condition numbers $\kappa\approx 163,146,132$. Bottom-right: $r_{\rm loc}(t)$ with different condition numbers. Here $d=1,000$ and $r_{0}=10$.
First, we show in the top-left plot in Figure 1 that the $|P_{t}(i,j)|$ is indeed exponentially decaying with $|i-j|$. Here we take a snapshot of the precision matrix at $t=0.1039$, which is the time with maximal effective localization radius (see bottom-left plot in Figure 1). We note that the precise exponential decay is not chosen artificially, and any snapshot will yield similar results.
@@
From Figure 1, we can see that the effective localization radius $r_{\rm loc}(t)$ first increases with $t$, and then decreases to $1$ when $t$ is large. This is due to the fact that $P_{t}$ can be regarded as an interpolation between $P_{0}$ and $P_{\infty}=I$. Note this is consistent with the theoretical prediction in 2.1, where the bound of $\|\nabla_{ij}^{2}\log p_{t}\|$ first increases with $t$ and then decreases to 00. Next, we can see that the effective localization radius $r_{\rm loc}(t)$ is almost independent of the dimension $d$, consistent with our motivation that the locality structure is approximately preserved with dimension independent radius. We can also see that the effective localization radius $r_{\rm loc}(t)$ is almost linear in the bandwidth $r_{0}$, and increases with the condition number $\kappa$.
@@ 4.1.2 Balance of localization error and statistical error
$X_{1}\sim\mathsf{N}(0,1),\quad X_{n+1}=\alpha_{h}X_{n}+\sigma_{h}\xi_{n},\quad\xi_{n}\sim\mathsf{N}(0,1),$
@@
We will use (4.5) to sample the target distribution with different localization radii $r$, and compare it to the reference sampling process (4.4). Although the localized score $\widehat{s}^{{\rm loc},r}(t,x)=-\widehat{P}_{t}^{{\rm loc},r}x$ in (4.5) is not the minimzer of ^$\widehat{\mathcal{L}}_{N}(s_{\theta})$ (3.6), it is very close to the minimizer, and it still yields a good approximation.
@@ 4.2 Cox-Ingersoll-Ross model
$\frac{X(t)}{c(t)}\sim H(t),\quad c(t)=\frac{\sigma^{2}}{8a}(1-e^{-2at}),$ (4.8)
where $H(t)$ is a noncentral $\chi$-squared distribution with $8ab/\sigma^{2}$ degrees of freedom and noncentrality parameter $c(t)^{-1}e^{-2at}X(0)$.
@@
For the diffusion model we choose a linear variance schedule with $\beta(t)=(\beta_{T}-\beta_{0})t/T+\beta_{0}$ with $T=0.05$, $\beta_{T}=0.5$ and $\beta_{0}=0.0001$, and where we sample the diffusion time $t\in[0,T]$ in steps of $0.001$ diffusion time units. The discount factor is given by $\alpha(t)=1-\beta(t)$. The score is estimated from $5,000$ randomly selected training points, differing in their uniformly sampled diffusion times and initial training sample. To learn the score function we employ a neural network with $3$ hidden layers of sizes $22r+$ 2, $6$ and $3$, respectively, with an input dimension of $22r+$ 2 coming from the localized states of dimension $12r+$ 1 and the diffusion time. The weights of the neural network are determined by minimizing the MSE error using an Adams optimizer with a learning rate $\eta=0.00005$.
@@
Figure 3: Comparison of the data obtained from the original CIR model (4.7) and from the diffusion model for localization radii $r=0$ (left), $r=2$ (middle) and $r=20$ (right). Top: Empirical histogram. Bottom: Auto-correlation function, averaged over all $2,500$ samples. The dashed lines mark deviations of the sample mean that are $1$ standard deviation away. The light grey lines show the individual auto-correlation functions of the generated data.
For the training we estimate the score function at entry $i$ for $i=2r+1,\ldots,N-2r-1$ from the localized state $(x_{r})_{i}=[x_{i-r},\ldots,x_{i},\ldots,x_{i+r}]\in\mathbb{R}^{2r+1}$. Due to stationarity of the process, each component of the score function $s_{i}((x_{r})_{i})$ will be the same except the boundaries, i.e. $i\leq r$ or $i\geq d-r$. This allows us to train a single score function which takes a $(2r+2)$-dimensional input ($12r+$ 1 for the localized state and $1$ for the diffusion time) to generate a $1$-dimensional output of the score function at location $r<i<d-r$. To deal with the boundaries of the time series for $i=1,\dots,r$ and $i=N-r,\ldots,N$, we pad with the time series $x$, reflected around $i$. During the training process we have employed independent noise for each localized region. We have checked that the results do not change if the noise in the diffusion model is kept constant for each local input or if varied when cycling through the localized regions.
@@ 5 Conclusions
Acknowledgements
The work of SR has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294. GAG acknowledges funding from the Australian Research Council, grant DP220100931. GAG thanks Yuguang Hu and Xiyu Wang for valuable discussions on the implementation of diffusion models. The work of SL is partially supported by NUS Overseas Research Immersion Award (ORIA). The work of XTT is supported by Singapore MOE grant A-8002956-$-0$
@@ Appendix A Proofs in Section 2
[…]
@@
[…]
@@
[…]
@@
$-\Delta u_{f}(x)-\nabla\log p(x)\cdot\nabla u_{f}(x)=f(x_{i}).$
@@
[…]
@@
$-\Delta_{p}u_{f}(x):=-\Delta u_{f}(x)-\nabla\log p(x)\cdot\nabla u_{f}(x)=f(x_{i})-\mathbb{E}_{p(x)}[f(x_{i})].$ (A.3)
@@
See also [13] for a detailed proof. Differentiating w.r.t $x_{j}$ gives
@@
Here $\nabla_{j}X_{t,i}^{x}$ is the partial derivative w.r.t $x_{j}$ of the sample path. Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{j}X_{t,i}^{x}$. Since $f$ is Lipschitz, we obtain
@@
[…]
@@ Appendix B Proofs in Section 3
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3.3.2 A Complete Error Analysis
  • (Boundedness) $p_{0}$ is supported on $M,M]^{d}$, and its density is upper and lower bounded by some constants $C_{p},C_{p}^{-1}$ respectively.
@@
[33] only considers the standard domain $1,1]^{d}$. It can be simply extended to $M,M]^{d}$ by scaling argument. Denote $p^{M}:=M^{d}p_{0}(M\cdot)$, then $p^{M}$ is supported on $1,1]^{d}$ and satisfies the same regularity conditions. Note the scaling only affects the radius $R$ of the Besov space, and does not change the scaling of the sample complexity.
@@
$\mathsf{L}^{j}=\mathcal{O}(\log^{4}n_{j}),\quad\left\|\mathsf{W}^{j}\right\|_{\infty}=\mathcal{O}(n_{j}\log^{6}n_{j}),\quad\mathsf{S}^{j}=\mathcal{O}(n_{j}\log^{8}n_{j}),\quad\mathsf{B}^{j}=n_{j}^{\mathcal{O}(\log\log n_{j})}.$ (3.24)
@@
$\mathrm{e}^{-cr^{*}}+N^{-\frac{2\gamma}{d_{\rm eff}^{*}+2\gamma}}\ll N^{-\frac{2\gamma}{d+2\gamma}}.$ (3.26)
@@
$\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{W}_{2}^{2}(p,\widehat{q}_{T-{\underline{t}}})]\lesssim\mathsf{W}_{2}^{2}(p,p_{\underline{t}})+\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{W}_{2}^{2}(p_{\underline{t}},\widehat{q}_{T-{\underline{t}}})]\lesssim dN^{-k}+\mathbb{E}_{\{X^{(i)}\}_{i=1}^{N}}[\mathsf{KL}(p_{\underline{t}}\|\widehat{q}_{T-{\underline{t}}})].$ (3.27)
@@ 4.1.1 Approximate locality
$P_{t}:=-\nabla^{2}\log p_{t}=(\alpha_{t}^{2}C_{0}+\sigma_{t}^{2}I)^{-1}.$ (4.3)
@@
Figure 1: Top-left: The precision matrix $P_{t}$ at $t=0.1039$, plotted in $\log|P_{t}|$ scale. We can see precise exponential decay of $P_{t}(i,j)$ in |ij||i-j|. The rest plots are the localization radius $r_{\rm loc}(t)$ (4.4) under different problem dimension $d$, precision matrix bandwidth $r_{0}$ and condition number $\kappa$. Top-right: $r_{\rm loc}(t)$ with different dimensions. Here $r_{0}=10$ and the condition numbers are similar ($\kappa\approx 193,191,197$). Bottom-left: $r_{\rm loc}(t)$ with different bandwidths. Here $d=1,000$ and condition numbers $\kappa\approx 163,146,132$. Bottom-right: $r_{\rm loc}(t)$ with different condition numbers. Here $d=1,000$ and $r_{0}=10$.
First, we show in the top-left plot in Figure 1 that the $|P_{t}(i,j)|$ is indeed exponentially decaying with |ij||i-j|. Here we take a snapshot of the precision matrix at $t=0.1039$, which is the time with maximal effective localization radius (see bottom-left plot in Figure 1). We note that the precise exponential decay is not chosen artificially, and any snapshot will yield similar results.
@@
From Figure 1, we can see that the effective localization radius $r_{\rm loc}(t)$ first increases with $t$, and then decreases to $1$ when $t$ is large. This is due to the fact that $P_{t}$ can be regarded as an interpolation between $P_{0}$ and $P_{\infty}=I$. Note this is consistent with the theoretical prediction in 2.1, where the bound of $\|\nabla_{ij}^{2}\log p_{t}\|$ first increases with $t$ and then decreases to 0. Next, we can see that the effective localization radius $r_{\rm loc}(t)$ is almost independent of the dimension $d$, consistent with our motivation that the locality structure is approximately preserved with dimension independent radius. We can also see that the effective localization radius $r_{\rm loc}(t)$ is almost linear in the bandwidth $r_{0}$, and increases with the condition number $\kappa$.
@@ 4.1.2 Balance of localization error and statistical error
$X_{1}\sim\mathsf{N}(0,1),\quad X_{n+1}=\alpha_{h}X_{n}+\sigma_{h}\xi_{n},\quad\xi_{n}\sim\mathsf{N}(0,1),$ (4.5)
@@
We will use (4.9) to sample the target distribution with different localization radii $r$, and compare it to the reference sampling process (4.8). Although the localized score $\widehat{s}^{{\rm loc},r}(t,x)=-\widehat{P}_{t}^{{\rm loc},r}x$ in (4.9) is not the minimzer of $\widehat{\mathcal{L}}_{N}(s_{\theta})$ (3.6), it is very close to the minimizer, and it still yields a good approximation.
@@ 4.2 Cox-Ingersoll-Ross model
$\frac{X(t)}{c(t)}\sim H(t),\quad c(t)=\frac{\sigma^{2}}{8a}(1-\mathrm{e}^{-2at}),$ (4.12)
where $H(t)$ is a noncentral $\chi$-squared distribution with $8ab/\sigma^{2}$ degrees of freedom and noncentrality parameter $c(t)^{-1}\mathrm{e}^{-2at}X(0)$.
@@
For the diffusion model we choose a linear variance schedule with $\beta(t)=(\beta_{T}-\beta_{0})t/T+\beta_{0}$ with $T=0.05$, $\beta_{T}=0.5$ and $\beta_{0}=0.0001$, and where we sample the diffusion time $t\in[0,T]$ in steps of $0.001$ diffusion time units. The discount factor is given by $\alpha(t)=1-\beta(t)$. The score is estimated from $5,000$ randomly selected training points, differing in their uniformly sampled diffusion times and initial training sample. To learn the score function we employ a neural network with $3$ hidden layers of sizes $2r+2$, 66 and $3$, respectively, with an input dimension of $2r+2$ coming from the localized states of dimension $2r+1$ and the diffusion time. The weights of the neural network are determined by minimizing the MSE error using an Adams optimizer with a learning rate $\eta=0.00005$.
@@
Figure 3: Comparison of the data obtained from the original CIR model (4.11) and from the diffusion model for localization radii $r=0$ (left), $r=2$ (middle) and $r=20$ (right). Top: Empirical histogram. Bottom: Auto-correlation function, averaged over all $2,500$ samples. The dashed lines mark deviations of the sample mean that are 11 standard deviation away. The light grey lines show the individual auto-correlation functions of the generated data.
For the training we estimate the score function at entry $i$ for $i=2r+1,\ldots,N-2r-1$ from the localized state $(x_{r})_{i}=[x_{i-r},\ldots,x_{i},\ldots,x_{i+r}]\in\mathbb{R}^{2r+1}$. Due to stationarity of the process, each component of the score function $s_{i}((x_{r})_{i})$ will be the same except the boundaries, i.e. $i\leq r$ or $i\geq d-r$. This allows us to train a single score function which takes a $(2r+2)$-dimensional input ($2r+1$ for the localized state and $1$ for the diffusion time) to generate a $1$-dimensional output of the score function at location r<i<drr<i<d-r. To deal with the boundaries of the time series for $i=1,\dots,r$ and $i=N-r,\ldots,N$, we pad with the time series $x$, reflected around $i$. During the training process we have employed independent noise for each localized region. We have checked that the results do not change if the noise in the diffusion model is kept constant for each local input or if varied when cycling through the localized regions.
@@ Appendix A Proofs in Section 2
$=$~ $\frac{1}{p_{t}(x_{t})}\int\left(-\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}^{2}}\right)\left(-\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}^{2}}\right)^{\rm T}\mathsf{N}(x_{t};\alpha_{t}x_{0},\sigma_{t}^{2}I)p_{0}(x_{0})\mathrm{d}x_{0}$
@@
$=$~ $\sigma_{t}^{-4}\mathbb{E}_{p_{0|t}(x_{0}|x_{t})}\left(x_{t}-\alpha_{t}x_{0}\right)\left(x_{t}-\alpha_{t}x_{0}\right)^{\rm T}$
@@
$=$~ $\sigma_{t}^{-4}\mathsf{Cov}_{p_{0|t}(x_{0}|x_{t})}\left(x_{t}-\alpha_{t}x_{0},x_{t}-\alpha_{t}x_{0}\right)$
$=$~ $\alpha_{t}^{2}\sigma_{t}^{-4}\mathsf{Cov}_{p_{0|t}(x_{0}|x_{t})}\left(x_{0},x_{0}\right),$
@@
$\Delta u_{f}(x)-\nabla\log p(x)\cdot\nabla u_{f}(x)=f(x_{i}).$
@@
$\leq$~ $\int\left\|\nabla_{j}u_{f}(x)\right\|\left\|\nabla g(x_{j})\right\|p(x)\mathrm{d}x\leq\frac{1}{m}\left(1-\frac{m}{M}\right)^{\mathsf{d}_{G}(i,j)}\left|f\right|_{\rm Lip}\left|g\right|_{\rm Lip}.$
@@
$\Delta_{p}u_{f}(x):=-\Delta u_{f}(x)-\nabla\log p(x)\cdot\nabla u_{f}(x)=f(x_{i})-\mathbb{E}_{p(x)}[f(x_{i})].$ (A.3)
@@
See also [13] for a detailed proof. Differentiating w.r.t. $x_{j}$ gives
@@
Here $\nabla_{j}X_{t,i}^{x}$ is the partial derivative w.r.t. $x_{j}$ of the sample path. Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{j}X_{t,i}^{x}$. Since $f$ is Lipschitz, we obtain
@@
$\leq$~ $\left|f\right|_{\rm Lip}\int_{0}^{\infty}\mathrm{e}^{-Mt}\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\frac{t^{k}(M-m)^{k}}{k!}\mathrm{d}t$
$=$~ $\left|f\right|_{\rm Lip}\frac{1}{M}\sum_{k=\mathsf{d}_{G}(i,j)}^{\infty}\left(1-\frac{m}{M}\right)^{k}=\frac{1}{m}\left(1-\frac{m}{M}\right)^{\mathsf{d}_{G}(i,j)}\left|f\right|_{\rm Lip}.$
@@ Appendix B Proofs in Section 3
$=$~ $\mathsf{KL}(p_{T}\|\mathsf{N}(0,I))+\int_{\underline{t}}^{T}\mathbb{E}_{x_{t}\sim p_{t}}\left[\left\|\widehat{s}(x_{t},t)-s(x_{t},t)\right\|^{2}\right]\mathrm{d}t.$
@@
$=$~ $\sum_{i:\mathsf{d}_{G}(i,j)>r}\mathbb{E}_{x^{\prime}\sim p_{t}}\left[\|\nabla_{ij}^{2}\log p_{t}(x^{\prime})\|_{\rm F}^{2}\Big|x_{\mathcal{N}_{j}^{r}}^{\prime}=x_{\mathcal{N}_{j}^{r}}\right]$
$\leq$~ $d_{j}\sum_{i:\mathsf{d}_{G}(i,j)>r}\frac{\alpha_{t}^{4}}{\sigma_{t}^{4}\left(m\sigma_{t}^{2}+\alpha_{t}^{2}\right)^{2}}\left(1-\frac{m\sigma_{t}^{2}+\alpha_{t}^{2}}{M\sigma_{t}^{2}+\alpha_{t}^{2}}\right)^{2\mathsf{d}_{G}(i,j)}.$
@@
$\leq$~ $\int_{0}^{T}\left[d_{j}\sum_{i:\mathsf{d}_{G}(i,j)>r}\left(m+\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}\right)^{-1}\frac{\alpha_{t}^{4}}{\sigma_{t}^{4}\left(m\sigma_{t}^{2}+\alpha_{t}^{2}\right)^{2}}\left(1-\frac{m\sigma_{t}^{2}+\alpha_{t}^{2}}{M\sigma_{t}^{2}+\alpha_{t}^{2}}\right)^{2\mathsf{d}_{G}(i,j)}\right]\mathrm{d}t$
$\leq$~ $d_{j}\sum_{k=r+1}^{\infty}|\{i:\mathsf{d}_{G}(i,j)=k\}|\int_{0}^{\infty}\frac{\alpha_{t}^{4}}{\sigma_{t}^{2}\left(m\sigma_{t}^{2}+\alpha_{t}^{2}\right)^{3}}\left(1-\frac{m\sigma_{t}^{2}+\alpha_{t}^{2}}{M\sigma_{t}^{2}+\alpha_{t}^{2}}\right)^{2k}\mathrm{d}t$
$\leq$~ $d_{j}\max\{1,m^{-1}\}\log\kappa\sum_{k=r+1}^{\infty}|\{i:\mathsf{d}_{G}(i,j)=k\}|(1-\kappa^{-1})^{2k}.$
@@
$=$~ $\sum_{k=r+1}^{\infty}|\mathcal{N}_{j}^{k}|\left[(1-\kappa^{-1})^{2k}-(1-\kappa^{-1})^{2(k+1)}\right]-|\mathcal{N}_{j}^{r}|(1-\kappa^{-1})^{2(r+1)}$
$\leq$~ $S\kappa^{-1}(2-\kappa^{-1})\sum_{k=r+1}^{\infty}k^{\nu}(1-\kappa^{-1})^{2k}\leq 2S\kappa^{-1}(1-\kappa^{-1})^{2r}\sum_{k=1}^{\infty}(k+r)^{\nu}(1-\kappa^{-1})^{2k}.$
@@
$\leq$~ $d_{j}\max\{1,m^{-1}\}\log\kappa\cdot 2S\kappa^{-1}(1-\kappa^{-1})^{2r}\cdot(r+1)^{\nu}\nu!(1-\kappa^{-1})^{2}\kappa^{2(\nu+1)}$
$=$~ $Cd_{j}(r+1)^{\nu}(1-\kappa^{-1})^{2(r+1)}.$
@@
$=$~ $\mathbb{E}_{x_{\mathcal{N}_{j}^{r}}\sim p_{t}}\left[\mathbb{E}_{x^{\prime}\sim p_{t}}\left[\|u_{\theta,j}(x_{\mathcal{N}_{j}^{r}},t)-u_{j}^{*}(x_{\mathcal{N}_{j}^{r}},t)+u_{j}^{*}(x_{\mathcal{N}_{j}^{r}},t)-s_{j}(x^{\prime},t)\|^{2}\Big|x_{\mathcal{N}_{j}^{r}}^{\prime}=x_{\mathcal{N}_{j}^{r}}\right]\right]$
$=$~ $\mathbb{E}_{x_{\mathcal{N}_{j}^{r}}\sim p_{t}}\left[\|u_{\theta,j}(x_{\mathcal{N}_{j}^{r}},t)-u_{j}^{*}(x_{\mathcal{N}_{j}^{r}},t)\|^{2}\right]$
@@
$=$~ $\int_{0}^{\infty}\frac{y}{\left(\kappa^{-1}+y\right)^{2}}\left(\frac{1}{1+y}\right)^{2k}\mathrm{d}y\leq\int_{0}^{\infty}\frac{y}{\left(\kappa^{-1}+y\right)^{2}}\left(\frac{1}{1+y}\right)^{2}\mathrm{d}y$
$<$~ $\int_{0}^{\kappa^{-1}}\kappa^{2}y\mathrm{d}y+\int_{\kappa^{-1}}^{1}\frac{\mathrm{d}y}{y}+\int_{1}^{\infty}\frac{\mathrm{d}y}{y^{3}}=1+\log\kappa\leq 2\log\kappa.$
@@
$=$~ $\mathbb{E}_{\epsilon_{t}\sim\mathsf{N}(0,I)}\left[\left\|u_{\theta,j}(\alpha_{t}x_{0,\mathcal{N}_{j}^{r}}+\sigma_{t}\epsilon_{t,\mathcal{N}_{j}^{r}},t)+\sigma_{t}^{-1}\epsilon_{t,j}\right\|^{2}\right].$
@@
$=$~ $\mathbb{E}_{z\sim\pi(z)}\mathbb{E}_{y\sim\pi(y|z)}\left[\left\|s_{\theta}(y)\right\|^{2}-2(s_{\theta}(y))^{\rm T}\nabla_{y}\log\pi(y|z)+\left\|\nabla_{y}\log\pi(y|z)\right\|^{2}\right]$
$=$~ $\mathbb{E}_{z\sim\pi(z)}\mathbb{E}_{y\sim\pi(y|z)}\left[\left\|s_{\theta}(y)\right\|^{2}+2\text{tr}\left(\nabla s_{\theta}(y)\right)+\left\|\nabla_{y}\log\pi(y|z)\right\|^{2}\right]$
$=$~ $\mathbb{E}_{y\sim\pi(y)}\left[\left\|s_{\theta}(y)\right\|^{2}+2\text{tr}\left(\nabla s_{\theta}(y)\right)+\left\|\nabla_{y}\log\pi(y)\right\|^{2}\right]+{\rm const}$
$=$~ $\mathbb{E}_{y\sim\pi(y)}\left\|s_{\theta}(y)-\nabla_{y}\log\pi(y)\right\|^{2}+{\rm const}.$
@@
$=$~ $\frac{1}{p_{t}(x_{\mathcal{N}_{j}^{r}})}\int\nabla_{j}\log p_{t}(x_{t,\mathcal{N}_{j}^{r}},x_{t,\mathcal{N}_{j}^{r_{\bot}}})p_{t}(x_{t,\mathcal{N}_{j}^{r}},x_{t,\mathcal{N}_{j}^{r_{\bot}}})\mathrm{d}x_{t,\mathcal{N}_{j}^{r_{\bot}}}$
$=$~ $\dfrac{\int\nabla_{j}p_{t}(x_{t,\mathcal{N}_{j}^{r}},x_{t,\mathcal{N}_{j}^{r_{\bot}}})\mathrm{d}x_{t,\mathcal{N}_{j}^{r_{\bot}}}}{\int p_{t}(x_{t,\mathcal{N}_{j}^{r}},x_{t,\mathcal{N}_{j}^{r_{\bot}}})\mathrm{d}x_{t,\mathcal{N}_{j}^{r_{\bot}}}}.$
@@
$=$~ $\dfrac{\int\left(-\sigma_{t}^{-2}(x_{t,j}-\alpha_{t}x_{0,j})\right)\mathsf{N}(x_{t,\mathcal{N}_{j}^{r}};\alpha_{t}x_{0,\mathcal{N}_{j}^{r}},\sigma_{t}^{2}I)p_{0}(x_{0,\mathcal{N}_{j}^{r}})\mathrm{d}x_{0,\mathcal{N}_{j}^{r}}}{\int\mathsf{N}(x_{t,\mathcal{N}_{j}^{r}};\alpha_{t}x_{0,\mathcal{N}_{j}^{r}},\sigma_{t}^{2}I)p_{0}(x_{0,\mathcal{N}_{j}^{r}})\mathrm{d}x_{0,\mathcal{N}_{j}^{r}}}.$
@@
$=$~ $\sum_{j=1}^{b}\mathbb{E}_{x_{t}\sim p_{t}}\left[\left\|\widehat{s}_{j}(x_{t},t)-s_{j}^{*}(x_{t},t)\right\|^{2}\right]+\sum_{j=1}^{b}\mathbb{E}_{x_{t}\sim p_{t}}\left[\left\|s_{j}^{*}(x_{t},t)-s_{j}(x_{t},t)\right\|^{2}\right].$
@@
$=$~ $\mathrm{e}^{-2T}\mathsf{KL}(p_{0}\|\mathsf{N}(0,I))+\int_{\underline{t}}^{T}\mathbb{E}_{x_{t}\sim p_{t}}\left[\left\|s^{*}(x_{t},t)-s(x_{t},t)\right\|^{2}\right]\mathrm{d}t+\mathcal{R}$
$\leq$~ $\mathrm{e}^{-2T}\mathsf{KL}(p_{0}\|\mathsf{N}(0,I))+Cd(r+1)^{\nu}\mathrm{e}^{-c(r+1)}+\mathcal{R},$
@@
Acknowledgments
GAG thanks Yuguang Hu and Xiyu Wang for valuable discussions on the implementation of diffusion models.
Funding
The work of SR has been funded by Deutsche Forschungsgemeinschaft (DFG) - Project-ID 318763901 - SFB1294. GAG acknowledges funding from the Australian Research Council, grant DP220100931. The work of SL is partially supported by NUS Overseas Research Immersion Award (ORIA). The work of XTT is supported by Singapore MOE grant A-8002956-00-00.